Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, December 11, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 11, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/11/98 (Friday)

Segment 1 report for 12/11/98: Calibration was still in progress at Emidio Pump Station. A segment of pipe was placed for tie-in to the Stream Crossing #4 bore (EMP's 8.92 to 8.98). Welding of pipe for the California Aqueduct bore occurred (EMP 12.28). Bore operations continued at Stream Crossing #9 (EMP 15.35) and excavation of the bore pit was in progress at the Tejon Bench/I-5 bore (EMP 16.49). On Grapevine Rd. coating occurred from EMP 15.53 to 15.56. Shading continued to EMP 15.95 working north. Electrical work and building up of the berm around Grapevine Pump Station continued. On Tejon Bench pipe was placed and tied-in to EMP 18.08. On Lebec Rd. excavation continued to EMP 21.60. Conduit was placed and backfill occurred to EMP 21.51. Pipe is being pushed through the Cuddy Creek Tributary bore (EMP 25.47).

Segment 2B (Castaic to Sylmar): Report for 12-11-98. Backfill, fiber optic conduit, x-ray and coating crews have completed mainline work on segment 2B. Crews completed work to tie-in the ditch line at Weldon Cyn Rd at EMP 71.4. In addition, asphalt has been completed and a box was set at EMP 71.4. Further, crews cleaned debris out of gutters and removed shoring jacks and plates from the right of way. At the Santa Clara river, crews pulled 1300' of fiber optic duct and set two pull boxes at either side of the river. Meanwhile, fiber optic cable was pulled between EMP 69.99 and 71.4 and a ditch was completed across San Fernando Rd at Balboa Blvd to provide a drain pipe for the hydro testing of the segment.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 11 December 1998: 3A: Clean up work continued between Allen and Grandview today. NIGHT crews completed their grouting and insulation work on San Fernando and Hollywood Way Bridges. Fiber optic conduit was pulled through the bore at Tuxford Street today. Ditch crews began digging on the north side of Tuxford for the tie in to the bore. Welders worked on pipe strung near Tuxford. Coating crews spent the day working on the Pacoima Wash Bridge crossing. 3C: Crews finished placing the three pull boxes, set at Grapevine Creek, Sepulveda Boulevard, and at the aqueduct crossing. More lids were placed on top of the structure for the aqueduct crossing. Work on the high speed channel cover progressed. Crews worked late digging across the road a conduit tie in. Work progressed on the pipe at Hubbard Avenue and north.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page